The solution of y – x.y’ = a( y^2 + y’) [y’ represents the first derivative of y w.r.t. x]

Grade:12

1 Answers

Aditya Gupta
4 years ago
y – x.y’ = a( y^2 + y’)
y(1 – ay)= y’(x+a)= (x+a)dy/dx
or dx/(x+a)= dy/y(1 – ay)
now integrate both sides and the ques will be solved. for RHS you can use partial fractions to integrate
kindly approve :)